# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v1

@@ III-B Analysis
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation 3 and the attack scheme $f_{attack}$ being order-invariant as in equation 7, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}\lor 1}\right],$ (8)
@@
Unlike the original AdaDetect, which guarantees $\text{FDR}\leq\alpha$ in benign settings, Theorem 1 provides an upper-bound under adversarial perturbations. In our proofs, we start with decomposing the FDR into attacked and unattacked components, and our key technical innovations (Lemmas 1 and 2) say that the first term $\mathbb{E}[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}}\lor 1}]$ remains bounded by $\alpha$ even after perturbations, demonstrating that AdaDetect’s statistical control over unattacked samples is preserved.
@@
$U_{\setminus\mathcal{A}}$ $=(U_{1},\dots,U_{n+\tilde{m}_{0}}):=(Z_{1},\dots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}),$
@@
$(U_{k+1},\dots,U_{n+\tilde{m}_{0}})\mid V\cup\widetilde{U}_{\mathcal{A}}\overset{d}{=}(U_{\pi(k+1)},\dots,U_{\pi(n+\tilde{m}_{0})})\mid V\cup\widetilde{U}_{\mathcal{A}}$
@@
$(\tilde{s}(U_{k+1}),\ldots,\tilde{s}(U_{n+\tilde{m}_{0}}))\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
$\qquad\overset{d}{=}(\tilde{s}(U_{\pi(k+1)}),\ldots,\tilde{s}(U_{\pi(n+\tilde{m}_{0})}))\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
@@
Building on Lemma 1 and Lemma 2, the following result follows directly from [7, Theorem A.1 (iii) and (iv))] and we skip the proof.
@@
Under the Lemma 2 setting. Let $i\in\mathcal{H}_{0}$ be an unattacked null index and $S_{i}:=\tilde{s}(\widetilde{Z}_{i})$ for $i\in[1:m+n]$. Define
$\widetilde{W}_{i}\;=\;\{S_{k+1},\dots,S_{n},S_{n+i}\}\;\cup\;(S_{n+j}:j\neq i,j\in\mathcal{H}_{0})\;\cup\;(S_{n+j}:j\in\mathcal{H}_{1}).$ (10)
Then we have
  1. (i)
    The $p$-value ${\tilde{p}}_{i}$ is independent of $\widetilde{W}_{i}$.
  2. (ii)
    The quantity $(n-k+1){\tilde{p}}_{i}$ is uniformly distributed on the integers $\{1,\dots,n-k+1\}$.
@@ III-C Surrogate Decision-based Attack Scheme
  1. Data: Only test samples $\{Z_{j}\}_{j=n+1}^{m+n}$, but the attacker does not know which ones are nulls and non-nulls;
  2. Query: The attacker can query the user (who owns all the training and test data and the AdaDetect algorithm) to obtain the labels for all the testing data $\{Z_{j}\}_{j=n+1}^{m+n}$.
@@
  1. Step 1: Initial detection. The attacker queries the labels of the test data from the user. Upon request, the user applies AdaDetect to the full test set $\{Z_{n+i}\}_{i=1}^{m}$ at once, producing pseudo-labels $Y_{i}\in\{0,1\}$ where,
@@
  2. Step 2: Surrogate score function training. Assuming AdaDetect has reasonable detection power, we form the pseudo-labeled dataset $\mathcal{D}=\{(Z_{n+i},Y_{i})\}_{i=1}^{m}$ and train a surrogate score function
@@
  3. Step 3: Attack set selection and adversarial perturbation. Select a subset $\mathcal{A}$ from the unrejected test samples as the attack target, where $|\mathcal{A}|=m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$ with $\gamma\in(0,1]$ being the attack intensity parameter. For each $i\in\mathcal{A}$, compute
@@ IV Experiments
As explained in Section III-A, we will focus on two adversarial machine learning attacks: HSJA and Boundary attack. Each experiment is repeated $20$ times to calculate the mean and variance of the FDR and power. We denote the AdaDetect score function by $s(z)$ and the score function used for attack by g(z){g}(z). The estimated upper bound according to equation 8 is computed by $\alpha+m_{a}\cdot\frac{1}{20}\sum_{i=1}^{20}\left(\frac{1}{{\widetilde{R}}^{(i)}\lor 1}\right)$. This is computed for all of the rest of the experiments and they are listed in the tables.
IV-A Synthetic Data Experiments
We conduct four comprehensive experiments to evaluate our method’s performance across different data distributions and model configurations. All experiments use the following base parameters: training sample size $n=5000$, testing sample size $m=1000$, $k=4000$, true null data $m_{0}=900$, and significance level $\alpha=0.1$. We test our approach on three types of synthetic data distributions: independent Gaussian, non-Gaussian, and exchangeable Gaussian data. Due to space limitations, we defer the synthetic data generation and one more synthetic experiments in Appendix E.
Experiment 1: Varying attack size with random forest (RF) models.
@@ IV-A Synthetic Data Experiments
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original FDR $08\pm 0.03 0.0$$8\pm 0.04 0.0$$08\pm 0.04$
oracle ($m_{a}=50$) $0.36\pm 0.02$ $0.40\pm 0.05$ $0.38\pm 0.02$
surrogate ($m_{a}=50$) $0.34\pm 0.02$ $0.20\pm 0.05$ $0.37\pm 0.02$
estimated upper bound $0.43$ $0.42$ $0.41$
oracle ($m_{a}=200$) $0.67\pm 0.00$ $0.71\pm 0.01$ $0.69\pm 0.00$
surrogate ($m_{a}=200$) $0.67\pm 0.00$ $0.64\pm 0.01$ $0.67\pm 0.00$
estimated upper bound $0.80$ $0.78$ $0.77$
TABLE II: Experiment 1: Power + RF
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original power $0.96\pm 0.02$ $0.55\pm 0.06$ $1.00\pm 0.00$
oracle ($m_{a}=50$) $0.99\pm 0.01$ $0.87\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=50$) $0.96\pm 0.02$ $0.65\pm 0.05$ $1.00\pm 0.00$
oracle ($m_{a}=200$) $0.99\pm 0.01$ $0.98\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=200$) $0.96\pm 0.02$ $0.78\pm 0.05$ $1.00\pm 0.00$
IV-B Real-world Data Experiments
We evaluated the attack performance on four real-world datasets with diverse characteristics and application domains. All experiments use the same base parameters ($n$, $m$, $k$, $m_{0}$, $\alpha$) as the synthetic data experiments. Each experiment is repeated $20$ times to calculate the mean and variance of the FDR and power. We consider four popular datasets: Shuttle, Credit Card, KDDCup99, and Mammography (see data descriptions and one more experiment in Appendix E due to space limitations).
Experiment 2: Real-world data with RF models.
We evaluated both oracle and surrogate attack performance in the five real-world datasets using identical RF architectures for both score functions, with attack size $m_{a}=200$. This allows us to compare attack effectiveness across different real-world data characteristics. We compare the boundary attack with our default HSJA under both oracle and surrogate attack schemes, using identical RF architectures for both score functions with $m_{a}=200$. The result shows that our oracle and surrogate attack schemes and decision-based algorithms (HopSkipJump and Boundary) can significantly increase the FDR in comparison to the original FDR.
TABLE III: Experiment 2: FDR + RF
@@ IV-B Real-world Data Experiments
TABLE IV: Experiment 2: Power + RF
@@
We make two important observations. Firstly, compared to the other three datasets, the original power on Mammography is relatively low ($\sim 0.48$). As a consequence, the surrogate method learns the score function from less accurate labels, leading to a larger gap in FDR between the oracle vs. the surrogate. Similar phenomena also appear in the next real-world experiments, as well as the two in the Appendix E. Secondly, the power after attack often increases, since the attack targets data points near the decision boundary in practice, some of which belong to the alternative hypothesis and are thus more likely to be correctly rejected.
Experiment 3: Real-world data with mismatched configurations.
We apply four mismatched score function configurations (RF–NN, RF–RF) to the five real-world datasets with a fixed attack size of $m_{a}=200$, evaluating both the oracle and surrogate attack performance. This enables us to assess how different model combinations affect each attack type’s performance across various real-world scenarios. The results indicate that an attacker can employ a model different from the user’s and still inflate the FDR beyond the target level. Our experiments show that using the neural network configuration as the attacker’s model can substantially speed up the attack process. The results for RF-RF configuration is covered in Experiment $2$.
TABLE V: Experiment 3: FDR + RF-NN
@@ Appendix A Proof of Lemma 1
$\mathrm{P}\left(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}}\mid\widetilde{U}_{\mathcal{A}}=\tilde{u}_{\mathcal{A}},V=v\right)=\frac{\mathrm{P}(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}},\ f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}},\ V=v)}{\mathrm{P}(f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}},\ V=v)}.$ (14)
@@
$\mathrm{P}(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}},\ f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}\mid V=v)$ (15)
$=\mathrm{P}(U_{\setminus\mathcal{A}}^{\pi}\in\mathcal{U}_{\setminus\mathcal{A}},\ f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}\mid V=v).$ (16)
@@
Appendix B Appendix: Proof of Lemma 2
@@
$(S_{k+1},\dots,S_{n}),S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid\widetilde{U}_{\mathcal{A}},V,Q$
$\overset{d}{=}(S_{\pi(k+1)},\dots,S_{\pi(n)}),S_{\pi(n+i)}):i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid\widetilde{U}_{\mathcal{A}},V,Q.$
Now, we justify why conditioning on $(\widetilde{U}_{\mathcal{A}},V,Q)$ is equivalent to conditioning only on $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$. First, note that $Q=h(U,\widetilde{U}_{\mathcal{A}},V)$ is a deterministic function of $U$, $\widetilde{U}_{\mathcal{A}}$ and $V$. The score function $s$ depends on $\widetilde{U}_{\mathcal{A}}$ and $V$ only through this transformation $Q$.
Furthermore, those conditional scores are the results of a deterministic function of $(Q,\widetilde{U}_{\mathcal{A}},V)$. Therefore, the scores $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1}\cup\mathcal{A})$ are fully determined once $\widetilde{U}_{\mathcal{A}}$, $V$ and $Q$ are fixed, and vice versa.
@@
$(\tilde{s}(Z_{k+1}),\dots,\tilde{s}(Z_{n}),\tilde{s}(Z_{n+i}):i\in\mathcal{H}_{0}\setminus\mathcal{A})\text{ is exchangeable conditional on}$
$(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A}).$

### v2

@@ III-B Analysis
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation II-A and the attack scheme $f_{attack}$ being order-invariant as in equation 7, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (8)
@@
Unlike the original AdaDetect, which guarantees $\text{FDR}\leq\alpha$ in benign settings, Theorem 1 provides an upper-bound under adversarial perturbations. In our proofs, we start with decomposing the FDR into attacked and unattacked components, and our key technical innovations (Lemmas 1 and 2) say that the first term $\mathbb{E}[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}]$ remains bounded by $\alpha$ even after perturbations, demonstrating that AdaDetect’s control over unattacked samples is preserved.
@@
$U_{\setminus\mathcal{A}}$ $=(U_{1},\dots,U_{n+\tilde{m}_{0}}):=$
$\quad\quad\quad\quad(Z_{1},\dots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}),$
@@
$\begin{split}&(U_{k+1},\dots,U_{n+\tilde{m}_{0}})\mid V\cup\widetilde{U}_{\mathcal{A}}\overset{d}{=}\\ &\quad\quad\quad\quad\quad(U_{\pi(k+1)},\dots,U_{\pi(n+\tilde{m}_{0})})\mid V\cup\widetilde{U}_{\mathcal{A}}\end{split}$
@@
$(\tilde{s}(U_{k+1}),\ldots,\tilde{s}(U_{n+\tilde{m}_{0}}))\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup$
$(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})\overset{d}{=}(\tilde{s}(U_{\pi(k+1)}),\ldots,\tilde{s}(U_{\pi(n+\tilde{m}_{0})}))\mid$
$\quad\quad\quad\quad\quad(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
@@
Building on Lemma 1 and Lemma 2, the following result follows directly from Theorem A.1 (iii) and (iv) from [7] and we skip the proof.
@@
Under the Lemma 2 setting. Let $i\in\mathcal{H}_{0}$ be an unattacked null index and $S_{i}:=\tilde{s}(\widetilde{Z}_{i})$ for all $i$. Define
$\widetilde{W}_{i}\;=\;$ $\{S_{k+1},\dots,S_{n},S_{n+i}\}\;\cup\;$
$\quad(S_{n+j}:j\neq i,j\in\mathcal{H}_{0})\;\cup(S_{n+j}:j\in\mathcal{H}_{1}).$
Then we have,
(i) The $p$-value ${\tilde{p}}_{i}$ is independent of $\widetilde{W}_{i}$.
(ii) The quantity $(n-k+1){\tilde{p}}_{i}$ is uniformly distributed on the integers $\{1,\dots,n-k+1\}$.
@@ III-C Surrogate Decision-based Attack Scheme
Data: Only test samples $\{Z_{j}\}_{j=n+1}^{m+n}$, but the attacker does not know which ones are nulls and non-nulls;
Query: The attacker can query the user (who owns all the training and test data and the AdaDetect algorithm) to obtain the labels for all the testing data $\{Z_{j}\}_{j=n+1}^{m+n}$.
@@
Step 1: Initial detection. The attacker queries the labels of the test data from the user. Upon request, the user applies AdaDetect to the full test set $\{Z_{n+i}\}_{i=1}^{m}$ at once, producing pseudo-labels $Y_{i}\in\{0,1\}$ where,
@@
Step 2: Surrogate score function training. Assuming AdaDetect has reasonable detection power, we form the pseudo-labeled dataset $\newline \mathcal{D}=\{(Z_{n+i},Y_{i})\}_{i=1}^{m}$ and train a surrogate score function
@@
Step 3: Attack set selection and adversarial perturbation. Select a subset $\mathcal{A}$ from the unrejected test samples as the target, where $|\mathcal{A}|=m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$ with $\gamma\in(0,1]$ being the attack intensity parameter. For each $i\in\mathcal{A}$, compute
@@ IV Experiments
As explained in Section III-A, we consider both the Adadetect scheme and the one by Bates el al. [4], using two adversarial machine learning attacks: HSJA and Boundary attack.
IV-A Applying Surrogate Scheme on Adadetect
We evaluated the attack performance on three types of synthetic data distributions: independent Gaussian, non-Gaussian, and exchangeable Gaussian data, and four real-world datasets (Shuttle, Credit Card, KDDCup99, and Mammography) with diverse characteristics and application domains (see data descriptions in Appendix F). All experiments use the following base parameters: training sample size $n=5000$, testing sample size $m=1000$, $k=4000$, true null data $m_{0}=900$, and significance level $\alpha=0.1$. Each experiment is repeated $20$ times to calculate the mean and variance of the FDR and power.
Due to space limitations, we defer the synthetic data generation and experiment results to Appendix F, along with one more real-world data experiment.
Experiment 1: Real-world data with RF models. We evaluated both oracle and surrogate attack performance in the four real-world datasets using identical RF architectures for both score functions, with attack size $m_{a}=200$. This allows us to compare attack effectiveness across different real-world data characteristics. We compare the boundary attack with our default HSJA under both oracle and surrogate attack schemes, using identical RF architectures for both score functions with $m_{a}=200$. The result shows that our oracle and surrogate attack schemes and decision-based algorithms (HopSkipJump and Boundary) can significantly increase the FDR in comparison to the original FDR.
@@ IV-A Applying Surrogate Scheme on Adadetect
TABLE II: Experiment 1: Power + RF
@@
We make two important observations. Firstly, compared to the other three datasets, the original power on Mammography is relatively low ($\sim 0.48$). As a consequence, the surrogate method learns the score function from less accurate labels, leading to a larger gap in FDR between the oracle vs. the surrogate. Similar phenomena appear in the next real-world experiments, as well as the two in the Appendix F. Secondly, the power after attack often increases, since the attack targets data points near the decision boundary in practice, some of which belong to the alternative hypothesis and are thus more likely to be correctly rejected.
Experiment 2: Real-world data with mismatched configurations. We apply mismatched score function configurations (RF–NN) to the four real-world datasets with a fixed attack size of $m_{a}=200$, evaluating both the oracle and surrogate attack performance. This enables us to assess how different model combinations affect each attack type’s performance across various real-world scenarios. The results indicate that an attacker can employ a model different from the user’s and still inflate the FDR beyond the target level. Our experiments show that using the neural network configuration as the attacker’s model can substantially speed up the attack process. The results for RF-RF configuration is covered in Experiment $2$.
TABLE III: Experiment 2: FDR + RF-NN
@@
TABLE IV: Experiment 3 with Monte Carlo adjustment
Dataset Credit-card Shuttle Gaussian
original FDR $09\pm 0.02 0.0$$08\pm 0.04$ $0.10\pm 0.02$
surrogate+hop. $0.45\pm 0.08$ $0.58\pm 0.08$ $0.56\pm 0.09$
surrogate+bound. $0.41\pm 0.10$ $0.53\pm 0.07$ $0.69\pm 0.06$
TABLE V: Experiment 3 without Monte Carlo adjustment
Dataset Credit-card Shuttle Gaussian
original FDR $08\pm 0.01 0.0$$9\pm 0.02 0.0$$09\pm 0.03$
surrogate+hop. $0.45\pm 0.08$ $0.49\pm 0.10$ $0.58\pm 0.07$
surrogate+bound. $0.48\pm 0.07$ $0.54\pm 0.06$ $0.65\pm 0.09$
IV-B Adapting Our Surrogate Approach to Bates el al. [4] as in Section II-B
To demonstrate the generality of our attack framework, we evaluate the adversarial robustness of the one-class classifier-based method proposed by Bates et al. in [4]. This approach serves as a fundamental baseline. We utilize two real-world datasets and a synthetic Gaussian data distribution. The null (e.g., non-fraud) data is partitioned into a training set of size $n_{\text{train}}=4000$ and a calibration set of size $n_{\text{cal}}=1000$. The test set $\mathcal{D}_{\text{test}}$ consists of $m=2000$ samples, comprising $m_{0}=1800$ inliers (normal transactions) and $m_{1}=200$ outliers (e.g., fraud). The p-values are computed using one-class classifier-based scheme, and the BH procedure is applied for FDR control.
Surrogate Decision-based Attack Scheme. We implement the surrogate decision-based attack scheme described in Section III-C with specific adaptations:
(1) Target Selection: The attacker selects a set $\mathcal{A}$ of size $m_{a}=200$ from the unrejected test samples. To maximize attack efficiency, we select $200$ unrejected test samples with the smallest p-values, corresponding to samples naturally close to the decision boundary.
(2) Surrogate Training with Label Flipping: The attacker queries the one-class classifier-based scheme to obtain initial binary rejections. To train the surrogate model $g(z)$ (a Multi-Layer Perceptron with one hidden layer of 100 units), we employ a heuristic label-flipping strategy: $75\%$ of the rejected samples with the highest p-values are relabeled as inliers in the training set. This encourages the surrogate to learn a more restrictive decision boundary, as only the most extreme outliers retain their original labels. Since HSJA and Boundary attack are decision-based attack, this surrogate model with tighter decision boundary for outliers will encourage those two decision-based attack to be more aggressive. We remark that a more conservative boundary (e.g., using a $50\%$ threshold or no flip at all) often fails here because it results in a loose rejection region, making it difficult for the attack to successfully push a sample labeled inlier into the real outlier space. The sample may successfully move into this loose boundary for outliers, but it’s still not extreme enough for one-class classifier-based scheme to notice. While the optimal threshold may vary across different datasets, this parameter serves as a tuning knob: increasing the flipping percentage makes the HSJA and Boundary attack more aggressive, facilitating more effective adversarial generation.
(3) Adversarial Generation: We utilize HSJA and Boundary attack on the surrogate model $g(z)$ to generate adversarial perturbations for the target set $\mathcal{A}$. The attack is aiming to cross the decision boundary.
Experiment 3: Both real-world and synthetic data on Bates’s method. We evaluated the attack performance at significance levels $\alpha=0.1$ both with and without Monte Carlo adjustment (recall Remark 1). The results quantify whether our surrogate approach succeed at attacking the standard one-class classifier-based scheme.
@@ Appendix A Proof of Lemma 1
$\mathrm{P}\left(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}}\mid\widetilde{U}_{\mathcal{A}}=\tilde{u}_{\mathcal{A}},V=v\right)=\frac{\mathrm{P}(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}},\ V=v)}{\mathrm{P}(f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}},\ V=v)}.$
@@
$\mathrm{P}(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}|V=v)=\mathrm{P}(U_{\setminus\mathcal{A}}^{\pi}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}|V=v).$
@@
Appendix B Proof of Lemma 2
@@
$(S_{k+1},\dots,S_{n}),S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid\widetilde{U}_{\mathcal{A}},V,Q\overset{d}{=}(S_{\pi(k+1)},\dots,S_{\pi(n)}),S_{\pi(n+i)}):i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid\widetilde{U}_{\mathcal{A}},V,Q.$
Now, we justify why conditioning on $(\widetilde{U}_{\mathcal{A}},V,Q)$ is equivalent to conditioning only on $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$. First, note that $Q=h(U,\widetilde{U}_{\mathcal{A}},V)$ is a deterministic function of $U$, $\widetilde{U}_{\mathcal{A}}$ and $V$. The score function $s$ depends on $\widetilde{U}_{\mathcal{A}}$ and $V$ only through this transformation $Q$. Furthermore, those conditional scores are the results of a deterministic function of $(Q,\widetilde{U}_{\mathcal{A}},V)$. Therefore, the scores $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1}\cup\mathcal{A})$ are fully determined once $\widetilde{U}_{\mathcal{A}}$, $V$ and $Q$ are fixed, and vice versa.
@@
$(\tilde{s}(Z_{k+1}),\dots,\tilde{s}(Z_{n}),\tilde{s}(Z_{n+i}):i\in\mathcal{H}_{0}\setminus\mathcal{A})$ is exchangeable conditional on $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$.